OpenStudy (anonymous):

suppose y varies directly with x. if x is -7 when y is 3 , what is x when y is -5?

OpenStudy (amistre64):

y = k*x 3 = k * -7 -3/7 = k ................ -5 = (-3/7)x x = -5(-7)/3

OpenStudy (anonymous):

-15

OpenStudy (amistre64):

35/3

OpenStudy (amistre64):

11' 2/3 perhaps :)